OpenStudy (anonymous):

5 cos^4 (6x) Which identity do I use, and how do I work the problem?

OpenStudy (anonymous):

you need to simplify it?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

I was going to use cos2x = (1+cos(2x))/2

OpenStudy (anonymous):

but I am just having trouble with the operations

OpenStudy (anonymous):

i think you have written wrong you can use the identity\[\cos2x=2\cos ^{2}x-1\]

OpenStudy (anonymous):

So it would look like 5(2cos^2(6x)-1)^2 Is that right?

OpenStudy (anonymous):

from this you will get\[\frac{ 5 }{ 2 }*[ 1+\cos12x] \]

OpenStudy (anonymous):

do you get that

OpenStudy (anonymous):

in the above identity of cos2x, replace \[\cos ^{2}x\] with\[\cos ^{2}6x\] you get 1+cos2(6x)

OpenStudy (anonymous):

one mistake

OpenStudy (anonymous):

Not cos(12x) - 1? From using the identity cos2x = 2cos^2 (x) - 1

OpenStudy (anonymous):

see ther is \[\cos ^{2}x\] which is to be replaced by\[\cos ^{2}6x\] in the identity

OpenStudy (anonymous):

sorry and the correct ans. is

OpenStudy (anonymous):

\[5*(\frac{ 1+\cos12x }{ 2 })^{2}\]

OpenStudy (anonymous):

what is the thing which you dont got

OpenStudy (anonymous):

why there is a 2 in the denominator?

OpenStudy (anonymous):

I think we are talking about different identities now

OpenStudy (anonymous):

If we use cos^2 (x) = (1+cos2x)/2 Then we have 5((1+cos(12x))/2)^2

OpenStudy (anonymous):

I understand it like that

OpenStudy (anonymous):

I was going to use cos2x = (1+cos(2x))/2-------------this is that which you had written here cos2x in LHS is \[\cos ^{2}x\]

OpenStudy (anonymous):

Oh ok, I understand then. I thought you were using cos2x = 2cos^2 x - 1 That is what you wrote at the beginning. But I understand now. Thank you!

OpenStudy (anonymous):

yes what you have written is correct

OpenStudy (anonymous):

and then i FOIL those terms, distribute the 5 back in and badabing, I have my answer right?

OpenStudy (anonymous):

did the ans. match with that in ans. book

OpenStudy (anonymous):

The answer in the book is

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

@iambatman

OpenStudy (anonymous):

I have the answer, with the work

OpenStudy (anonymous):

@lall thank you!

OpenStudy (anonymous):

nice!